At what angle shall a curve be banked in order that a car with a velocity of 20.0 m/s will make the turn if the diameter of the path is 40.0 m?


1
Expert's answer
2021-11-29T11:44:58-0500

Let's apply the Newton's Second Law of Motion in projections on "x"- and "y"-axis:


"Nsin\theta=\dfrac{mv^2}{r},""Ncos\theta=mg."

Dividing the first equation by the second one, we get:


"tan\theta=\dfrac{v^2}{gr},""\theta=tan^{-1}(\dfrac{v^2}{gr}),""\theta=tan^{-1}(\dfrac{(20\ \dfrac{m}{s})^2}{9.8\ \dfrac{m}{s^2}\times20\ m})=63.9^{\circ}."